# A Time-Intensity Aware Pipeline for Generating Late-Stage Breast DCE-MRI using Generative Adversarial Models

### v1

@@ 3.2 Image normalization based on the Time-Intensity pattern
Image normalization plays an important role in image synthesis. Most related works use conventional normalization techniques such min-max and z-score normalization. However, normalizing pixel intensity based on single image statistics can distort diagnostic information because it varies temporally with other precontrast and postcontrast sequences. For this reason, we designed a normalization strategy to ensure that the information in the TI curve is preserved across image sequences. This technique is based on the Z-score standardization with respect to a different reference. For a given precontrast image $x_{p}$, both $\mu$ and $\sigma$ are computed as in the Z-score and the normalized image $x_{p_{TI}}$ is calculated. However, the early ($x_{e}$), real late ($y_{l}$) and generated late ($y_{l^{\prime}}$) images are standardized using the $\mu$ and $\sigma$ values from $x_{p}$. This will ensure that pixel values for all the postcontrast images will reflect the variation in terms of intensity with respect to the precontrast to effectively preserve the contrast enhancement pattern. We named this as Time-Intensity nomralization (TI-norm). To compute the TI-norm of $x_{p},x_{e},y_{l}$ and $y_{l^{\prime}}$, we use equations in 13
@@ 3.3 Quality evaluation based on Time-Intensity curve
Since the quality of our proposed approach depends on the ability of replicate the expected attributes of the time-intensity curve, we proposed two metrics based on its behavior, besides conventional pixel-based metrics. For a set of precontrast ($x_{p}$), early ($x_{e}$), late ($y_{l}$) and generated ($y_{l^{\prime}}$) responses, a real and a synthetic time-intensity curve is computed and three groups of metrics are estimated:
@@ 3.3.1 CA pattern score in annotated regions ($\mathcal{CP}_{s}$)
This metric aims to determine if the CA pattern (plateau, persistent and wash-out) in a region of the image with a small field of view is the same as expected. This approximation aligns with the medical practice, where the TI curve is computed in a single or small valley of pixels to analyze its behavior across multiple image sequences. For a set of regions in all image modalities ($Rx_{p}$, $Rx_{e}$, $Ry_{l}$ or $Rx_{p}$, $Rx_{e}$, $Ry_{l^{\prime}}$) the time intensity curve is computed. To determine the CA pattern for each of them, we calculate the percentage of enhancement ($\mathcal{E}_{ps}$) of the late response w.r.t. the early phase as shown in equation 16. Then, the type of TI curve is estimated following the standard rules shown in equation 15. Finally, the overall multi-class F1 score between them is computed and reported.
@@ 3.3.2 Average difference of $\mathcal{E}_{ps}$ in regions ($\mathcal{ED}$)
The purpose of this metric is to estimate the difference among both real and generated TI responses within the tissue under different conditions. In this case, the $\mathcal{E}_{ps}$ is computed in pairs of real and generated regions to determine how different the expected score is synthesized. This difference is calculated instead of the CA pattern given its clinical meaning, that is bounded to clinically relevant findings. Formally, given a set of annotated ROIs ($Rx_{p}$, $Rx_{e}$, $Ry_{l}$ or $Rx_{p}$, $Rx_{e}$, $Ry_{l^{\prime}}$) or a set of unannotated regions ($Px_{p}$, $Px_{e}$, $Py_{l}$ or $Px_{p}$, $Px_{e}$, $Py_{l^{\prime}}$) in all image modalities, the TI curve is calculated, as well as the percentage of enhancement ($\mathcal{E}_{ps}$). Then, the $L1$ norm is used to calculate the distance between the real and generated TI curves at late phase from the annotated ROIs ($\mathcal{ED}_{R}$) or unannotated regions ($\mathcal{ED}_{UR}$). The formulation of the score for both scores is shown in equation 16.
@@ 3.3.3 Image quality metrics
The purpose of this set of metrics is to evaluate the quality of the images in terms of spatial distribution. For this, we computed the Mean Absolute Error (MAE) to estimate the pixel-to-pixel change among real and generated sequences via spectral fidelity, the Structural Similarity Index Measure (SSIM) to estimate the structural quality of visual perception from three perspectives: correlation loss, luminance distortion, and contrast distortion, and the Peak Signal-to-Noise Ratio (PSNR) to determine the ratio between the maximum possible value (power) of the images and the power of distorting noise that affects their quality. Finally, to better understand the spatial differences in terms of intensity among the real and generated images, we visualized the absolute difference among each pixel intensity value as a heatmap.
4 Results and discussion
@@ 4.1 Dataset
To perform all experiments we used the Duke dataset [22]. This dataset includes 922 preoperative MRI imaging sequences of invasive breast cancer patients’ (precontrast and 3 or 4 postcontrast). The patients were between 21 and 89 years old and comprised seven racial and ethnic groups. Tumor size, nuclear grade, and hormone receptor status varied per patient. Ten scanner manufacturers were used to capture MRI images, and patients received three kinds of contrast agents. These constraints maintain heterogeneous parameters and prevent the data from being limited to one MRI configuration. Additionally, expert radiologists annotated at least one region with lesions or tissue anomalies per patient.
Our framework was built using precontrast, the earliest postcontrast (2mins after CA administration), and the latest sequence (6mins after CA administration). In our experiments, 1.5T and 3T images were used independently to train models due to their different visual features in terms of contrast enhancement. The 1.5T and 3T subsets had 393 and 389 patients, respectively. Due to sparse tissue contrast agent response, full DICOM volumes were not used in both cases. A frame-wise selection was performed on the basis of the available annotated ROIs, plus previous and posterior images. Thus, for training and validation, $9,196$ and $1,043$ images were extracted from the 3T images, and $8,941$ and $1,012$ from 1.5T. Additionally, images were normalized using TI-norm, as described in section 3.2.
@@ 4.2 Experimental setup
Regarding the PAN model, we ran three tests using the same training configuration and compiled the findings. For the baseline, we adopted the original implementation consisting of 16 SC-PA blocks and convolutions with a kernel size of 40 for downsampling and 24 for upsampling. Additionally, the discriminator network consists of a $1\times 1$ Markovian discriminator [40] with convolutional blocks (Conv2d + InstanceNorm + LeakyReLU) following the architecture as in [41]. Besides, using the same configuration, we trained the model including either CELoss (CE-PAN) or TILoss (TI-PAN) and compiled results to quantify the impact of the losses on the models’ performance.
To perform a comparative analysis considering the CA behavior, we evaluated the quality of the generated images at three different levels: full image, annotated ROIs and unannotated regions. The analysis of the full image is based on conventional pixel-wise metrics to check their realism as other similar works. On the other hand, the evaluation of ROIs annotated is based on the quality of the CA response generated using the metrics described in Section . As the annotations correspond to large portions of tissue, the entire region was not considered. Instead, a small set of pixels was taken ($3\times 3$), where its center corresponds to the maximum intensity point within the ROI.
However, since expected behavior in annotated ROIs could bias the analysis in terms of variability, we also analyzed the generated CA response in tissue regions without annotated clinical findings. For this, regions within the tissue portion of the organs were manually selected in $x_{e}$, and their locations were projected in the rest of the modalities to compile $Px_{p}$, $Px_{e}$, $Py_{l}$, $Py_{l^{\prime}}$. From each region, the TI curve and the increase in enhancement are computed as described in section 4.2. Additionally, to increase the number of samples to analyze, 3 unannotated regions are extracted from each image in the 3T test set, to complete $3,129$ unannotated regions. At this point, the type of TI curve was not considered, as this characterization is usually assigned to lesions or clinically relevant findings. Instead, we expect no significant difference between the generated and expected increase in the enhancement.
@@ 4.3 Clinical Performance based on TI curve metrics
Here, we evaluate the accuracy of of the generated images for clinical applications. To achieve this, we used our proposed pipeline to predict the late response images given the early response. Then, the diagnostic value and performance in terms of the TI curve for the annotated ROIs and unannotated regions is analyzed using the aforementioned metrics. Figure 3 a) shows the results for the model predicting the time-intensity response using TI-PAN with the 3T images, where patients are splitted according to their real CA pattern (persistent, plateau, and wash-out). The results show that the estimated CA pattern score for the annotated ROIs is $0.904$ and the expected TI response matches the real response for most patients, with small variations among them.
Additionally, in the most prominent error cases, the proposed pipeline avoids the over-estimation of the late CA response, which is important to avert false positive and false negative cases. Besides, in an in-depth analysis, only 5 cases were miss-estimated for the CA pattern. However, in all these cases the average difference between the real and generated enhancement was ±$\pm 2.5\%$, and the real enhancement was nearby the threshold that determines the type of TI curve (±$\sim\pm 10\%$). Also, all cases were either real persistent or wash-out that were estimated as plateau, which indicates that they are not conclusive and require further examination. These findings are relevant for our purpose since they demonstrates the ability of the model to replicate the expected response in the TI curve for the diagnosis reliably, while preserving the most important information for the clinical analysis and avoiding false positive and false negative cases.
@@
Finally, we aim to determine if there is a significant difference in the CA response between the real and generated group of regions according to their enhancement increase by using the non-parametric Wilcoxon signed rank test, where we chose a significance value of $0.05$. In the test, if the $p$-value rejects the null hypothesis, that means, there is a significant difference between the real and generated CA responses. Otherwise, if the $p$-value fails the test, it means that there is no evidence to support a difference between real and generated CA responses. For the annotated ROIs reported in the figure a), it can be observed that the $p$-values for the persistent, plateau and wash-out TI curves are not able to reject the null hypothesis in any case, which supports that the model is able to generate the CA response following the same real pattern with no significant difference.
However, the analysis on the real annotated ROIs alone exhibits a limited number of samples and cases to analyze, which might bias the evaluation of the quality of the synthesized CA response in terms of generalization. For this reason, we extended the evaluation using unannotated regions within the tissue zones to evidence the changes. Unlike the previous analysis, the determination of the kind of TI curve is not made at this stage due to its clinical significance, which is only done for relevant findings. Instead, the analysis is performed in terms of the expected increase in enhancement to avoid misinterpretations. Plot b) in figure 3 shows the distribution of the enhancement variation for the unannotated real and generated regions and it’s level of statistical significance using the aforementioned test. It can be observed that our pipeline is able to replicate the distribution of the increase in enhancement efficiently, with small differences around the mean. This is also demonstrated in the $p$-value obtained through the test, which indicates no evidence to support a significant difference in the increase in enhancement. Also, it worth notice that outliers for both distributions are located below $-0.25\%$ increase and over $+0.60\%$ increase. These thresholds are important since they indicate that the most susceptible cases of analysis correspond to regions where higher or lower enhancement increase must be analyzed by clinicians, which correlates correctly with the characterization of wash-out and persistent TI curves. For the within-the-distribution the cases, the results show that our pipeline is able to replicate the expected increase in enhancement with small variations.
@@ 4.4 Impact of image normalization in the TI-Curve and CA pattern
At this point, we trained TI-PAN using Min-Max, Z-score and TI-norm and generated the synthetic images. In figure 4, first row of a) shows the generated images from all models where no major visual difference is evident. However, to assess the impact of the TI-norm in the diagnostic quality of the images we computed the time-intensity curve of each annotated set of samples $Rx_{p},Rx_{e}$ $Ry_{l}$ and $Ry_{l^{\prime}}$. Using the RadiAnt software [42], we computed the TI curve for the images without any normalization and after normalization. The figure 4 displays time intensity plots in the second row, where it can be seen that the TI curve is destroyed when conventional Min-Max and Z-score are used, in comparison to the non-normalized one. Nevertheless, the TI-PAN model still tries to mimic the given behavior when using Min-Max or Z-score with certain similarity. This is a critical issue for the diagnosis since the information of the curve itself is degraded significantly, and it distorts its interpretability afterwards despite the high visual correlation among real and generated late post-contrast. Nevertheless, our proposed time-intensity normalization effectively preserves the CA behaviour in the real and generated images, which approaches the generated TI curve to a real clinical interpretation.
@@
We also extended the evaluation by comparing the predicted contrast enhancement difference and the image quality in annotated and unannotated 3T images, and results are shown in b). Although in this case the Z-score norm obtains best performance in terms of image quality, it falls very short in terms of the expected estimation of the CA pattern, achieving suboptimal performance. Similar behavior is also obtained using Min-Max norm, where pixel error is lower but diagnostic quality remains suboptimal. In both cases, the TI-PAN model shows a decrease in performance and exhibits a critical limitation in the clinical scenario, despite using TI-Loss. In the case of TI-norm, it guarantees the prevalence of the TI curve to be leveraged by the TI-Loss during training. These features allow to obtain best scores in the estimation of the CA pattern while achieving similar image-based quality (25.4 in PSNR). This makes TI-norm an important component in generating high-quality diagnostic information in an image synthesis setting, which has not been reported before in the medical imaging synthesis field.
@@ 4.5 Quantitative and Qualitative evaluation of synthesized Images
An evaluation in the visual quality of all the generated images at spatial feature level was performed. Figure 5 shows some generated sample results for images in the test set using all of the evaluated models together. Additionally, as images in the test set had at least an annotated ROI, they have been also compared visually. Finally, the difference heatmaps are also shown. In the case of 1.5T images, the generated samples show enough realism in the visual domain for all of the models, with exception of certain tissue textures and details. This is coherent with the previous clinical result where narrow margin of difference was obtained in the visual domain. However, in case of the 3T images, a smaller difference in the real and generated is evidenced when using the TILoss, especially in the annotated region where the difference in the brighter zone is lower. Besides, table 1 also reports the MAE, SSIM and PSNR for all models at full image and ROI leves, where no single model obtains the best scores consistently. In contrast, no large margin is evidenced among models, and slight differences are found. This result is in line with other cutting-edge studies that have shown good results in creating useful images from post-contrast sequences [18, 11].

### v2

@@ 2.2 Image normalization based on the Time-Intensity pattern
Image normalization plays an important role in image synthesis. Most related works use conventional normalization techniques such min-max and z-score normalization. However, normalizing pixel intensity based on single-image statistics can distort diagnostic information because it varies temporally with other precontrast and post-contrast sequences. For this reason, we designed a normalization strategy to ensure that the information in the TI curve is preserved across image sequences. This technique is based on the Z-score standardization with respect to a different reference. For a given precontrast image $x_{p}$, both $\mu$ and $\sigma$ are computed as in the Z-score and the normalized image $x_{p_{TI}}$ is calculated. However, the early ($x_{e}$), real late ($y_{l}$) and generated late ($y_{l^{\prime}}$) images are standardized using the $\mu$ and $\sigma$ values from $x_{p}$. This will ensure that pixel values for all the post-contrast images will reflect the variation in terms of intensity with respect to the precontrast to effectively preserve the contrast enhancement pattern. We named this as Time-Intensity nomralization (TI-norm). To compute the TI-norm of $x_{p},x_{e},y_{l}$ and $y_{l^{\prime}}$, we use equations in 13
@@ 2.3 Quality evaluation based on Time-Intensity curve
Since the quality of our proposed approach depends on the ability to replicate the expected attributes of the time-intensity curve, we proposed two metrics based on its behavior, in addition to conventional pixel-based metrics. For a set of pre-contrast ($x_{p}$), early ($x_{e}$), late ($y_{l}$) and generated ($y_{l^{\prime}}$) responses, a real and a synthetic time-intensity curve are computed, and three groups of metrics are estimated:
@@ 2.3.1 CA pattern score in annotated regions ($\mathcal{CP}_{s}$)
This metric aims to determine if the CA pattern (plateau, persistent, and wash-out) in a region of the image with a small field of view is the same as expected. This approximation aligns with medical practice, where the TI curve is computed in a single or small valley of pixels to analyze its behavior across multiple image sequences. For a set of regions in all image modalities ($Rx_{p}$, $Rx_{e}$, $Ry_{l}$ or $Rx_{p}$, $Rx_{e}$, $Ry_{l^{\prime}}$) the time intensity curve is computed. To determine the CA pattern for each of them, we calculate the percentage of enhancement ($\mathcal{E}_{ps}$) of the late response with respect to the early phase as shown in equation 16. Then, the type of TI curve is estimated following the standard rules shown in equation 15. Finally, the overall multi-class F1 score between them is computed and reported.
@@ 2.3.2 Average difference of $\mathcal{E}_{ps}$ in regions ($\mathcal{ED}$)
The purpose of this metric is to estimate the difference between both real and generated TI responses within the tissue under different conditions. In this case, the $\mathcal{E}_{ps}$ is computed in pairs of real and generated regions to determine how different the expected score is synthesized. This difference is calculated instead of the CA pattern given its clinical significance, which is bound to clinically relevant findings. Formally, given a set of annotated ROIs ($Rx_{p}$, $Rx_{e}$, $Ry_{l}$ or $Rx_{p}$, $Rx_{e}$, $Ry_{l^{\prime}}$) or a set of unannotated regions ($Px_{p}$, $Px_{e}$, $Py_{l}$ or $Px_{p}$, $Px_{e}$, $Py_{l^{\prime}}$) in all image modalities, the TI curve is calculated, as well as the percentage of enhancement ($\mathcal{E}_{ps}$). Then, the $L1$ norm is used to calculate the distance between the real and generated TI curves at the late phase from the annotated ROIs ($\mathcal{ED}_{R}$) or unannotated regions ($\mathcal{ED}_{UR}$). The formulation of the score for both scores is shown in equation 16.
@@ 2.3.3 Image quality metrics
The purpose of this set of metrics is to evaluate the quality of the images in terms of spatial distribution. For this, we computed the Mean Absolute Error (MAE) to estimate the pixel-to-pixel change between real and generated sequences via spectral fidelity, the Structural Similarity Index Measure (SSIM) to estimate the structural quality of visual perception from three perspectives: correlation loss, luminance distortion, and contrast distortion, and the Peak Signal-to-Noise Ratio (PSNR) to determine the ratio between the maximum possible value (power) of the images and the power of distorting noise that affects their quality. Finally, to better understand the spatial differences in terms of intensity among the real and generated images, we visualized the absolute difference among each pixel intensity value as a heatmap.
3 Results
@@ 3.1 Dataset
To perform all experiments, we used the Duke dataset  [31]. This dataset includes 922 preoperative MRI imaging sequences of invasive breast cancer patients (precontrast and 3 or 4 postcontrast). The patients were between 21 and 89 years old and comprised seven racial and ethnic groups. Tumor size, nuclear grade, and hormone receptor status varied among patients. Ten scanner manufacturers were used to capture MRI images, and patients received three kinds of contrast agents. These constraints maintain heterogeneous parameters and prevent the data from being limited to a single MRI configuration. Additionally, expert radiologists annotated at least one region with lesions or tissue anomalies for each patient.
Our framework was built using pre-contrast, the earliest post-contrast (2 mins after CA administration), and the latest sequence (6 mins after CA administration). In our experiments, 1.5T and 3T images were used independently to train models due to their different visual features regarding contrast enhancement. The 1.5T and 3T subsets included 393 and 389 patients, respectively. Due to the sparse tissue contrast agent response, full DICOM volumes were not used in either case. A frame-wise selection was performed based on the available annotated ROIs, as well as previous and subsequent images. Thus, for training and validation, $9,196$ and $1,043$ images were extracted from the 3T images, and $8,941$ and $1,012$ from 1.5T. Additionally, images were normalized using TI-norm, as described in section 2.2.
@@ 3.2 Experimental setup
Regarding the PAN model, we ran three tests using the same training configuration and compiled the findings. For the baseline, we adopted the original implementation consisting of 16 SC-PA blocks and convolutions with a kernel size of 40 for downsampling and 24 for upsampling. Additionally, the discriminator network consists of a $1\times 1$ Markovian discriminator  [34] with convolutional blocks (Conv2d + InstanceNorm + LeakyReLU) following the architecture as in  [35]. Furthermore, using the same configuration, we trained the model including either CELoss (CE-PAN) or TILoss (TI-PAN) and compiled results to quantify the impact of the losses on the models’ performance.
To perform a comparative analysis considering the CA behavior, we evaluated the quality of the generated images at three different levels: full image, annotated ROIs, and unannotated regions. The analysis of the full image is based on conventional pixel-wise metrics to check their realism, as in other similar works. On the other hand, the evaluation of annotated ROIs is based on the quality of the CA response generated using the metrics described in Section . As the annotations correspond to large portions of tissue, the entire region was not considered. Instead, a small set of pixels was taken ($3\times 3$), where its center corresponds to the maximum intensity point within the ROI.
However, since the expected behavior in annotated ROIs could bias the analysis in terms of variability, we also analyzed the generated CA response in tissue regions without annotated clinical findings. For this, regions within the tissue portion of the organs were manually selected in $x_{e}$, and their locations were projected in the rest of the modalities to compile $Px_{p}$, $Px_{e}$, $Py_{l}$, $Py_{l^{\prime}}$. From each region, the TI curve and the increase in enhancement are computed as described in section 3.2. Additionally, to increase the number of samples to analyze, three(3) unannotated regions are extracted from each image in the 3T test set, to complete $3,129$ unannotated regions. At this point, the type of TI curve was not considered, as this characterization is usually assigned to lesions or clinically relevant findings. Instead, we expect no significant difference between the generated and expected increase in enhancement.
@@ 3.3 Clinical Performance based on TI curve metrics
Here, we evaluate the accuracy of the generated images for clinical applications. To achieve this, we used our proposed pipeline to predict the late response images given the early response. Then, the diagnostic value and performance in terms of the TI curve for the annotated ROIs and unannotated regions are analyzed using the aforementioned metrics. Figure 2 a) shows the results for the model predicting the time-intensity response using TI-PAN with the 3T images, where patients are split according to their real CA pattern (persistent, plateau, and washout). The results show that the estimated CA pattern score for the annotated ROIs is $0.904$ and the expected TI response matches the real response for most patients, with small variations among them.
Additionally, in the most prominent error cases, the proposed pipeline avoids the overestimation of the late CA response, which is important to avert false positive and false negative cases. Besides, in an in-depth analysis, only 5 cases were misestimated for the CA pattern. However, in all these cases, the average difference between the real and generated enhancement was ±$\pm 2.5\%$, and the real enhancement was near the threshold that determines the type of TI curve (±$\sim\pm 10\%$). Moreover, all cases were either real persistent or wash-out that were estimated as plateau, which indicates that they are not conclusive and require further examination. These findings are relevant for our purpose since they demonstrate the ability of the model to replicate the expected response in the TI curve for diagnosis reliably, while preserving the most important information for clinical analysis and avoiding false positive and false negative cases.
@@
Finally, we aim to determine if there is a significant difference in the CA response between the real and generated groups of regions, according to their enhancement increase, by using the non-parametric Wilcoxon signed-rank test, where we chose a significance value of $0.05$. In the test, if the $p$-value rejects the null hypothesis, that means there is a significant difference between the real and generated CA responses. Otherwise, if the $p$-value fails the test, it means that there is no evidence to support a difference between real and generated CA responses. For the annotated ROIs reported in figure a), it can be observed that the $p$-values for the persistent, plateau, and wash-out TI curves are not able to reject the null hypothesis in any case, which supports that the model is able to generate the CA response following the same real pattern with no significant difference.
However, the analysis on the real annotated ROIs alone exhibits a limited number of samples and cases to analyze, which might bias the evaluation of the quality of the synthesized CA response in terms of generalization. For this reason, we extended the evaluation using unannotated regions within the tissue zones to highlight the changes. Unlike the previous analysis, the determination of the kind of TI curve is not made at this stage due to its clinical significance, which is only addressed for relevant findings. Instead, the analysis is performed in terms of the expected increase in enhancement to avoid misinterpretations. Plot b) in figure 2 shows the distribution of the enhancement variation for the unannotated real and generated regions and its level of statistical significance using the aforementioned test. It can be observed that our pipeline is able to replicate the distribution of the increase in enhancement efficiently, with small differences around the mean. This is also demonstrated in the $p$-value obtained through the test, which indicates no evidence to support a significant difference in the increase in enhancement. Additionally, it is worth noting that outliers for both distributions are located below $-0.25\%$ increase and above $+0.60\%$ increase. These thresholds are important since they indicate that the most susceptible cases of analysis correspond to regions where higher or lower enhancement increases must be examined by clinicians, which correlates correctly with the characterization of wash-out and persistent TI curves. For the within-the-distribution cases, the results show that our pipeline is able to replicate the expected increase in enhancement with small variations.
@@ 3.4 Impact of image normalization in the TI-Curve and CA pattern
At this point, we trained TI-PAN using Min-Max, Z-score, and TI-norm, and generated the synthetic images. In Figure 3, first row of a) shows the generated images from all models, where no major visual difference is evident. However, to assess the impact of the TI-norm on the diagnostic quality of the images, we computed the time-intensity curve of each annotated set of samples $Rx_{p},Rx_{e}$ $Ry_{l}$ and $Ry_{l^{\prime}}$. Using RadiAnt software  [36], we computed the TI curve for the images without any normalization and after normalization. Figure 3 displays time-intensity plots in the second row, where it can be seen that the TI curve is altered when conventional Min-Max and Z-score are used, in comparison to the non-normalized one. Nevertheless, the TI-PAN model still tries to mimic the given behavior when using Min-Max or Z-score with certain similarity. This is a critical issue for diagnosis since the information of the curve itself is degraded significantly, and it distorts its interpretability afterwards, despite the high visual correlation among real and generated late post-contrast images. Nonetheless, our proposed time-intensity normalization effectively preserves the CA behavior in the real and generated images, which approaches the generated TI curve to a real clinical interpretation.
We also extended the evaluation by comparing the predicted contrast enhancement difference and the image quality in annotated and unannotated 3T images, and the results are shown in b). Although, in this case, the Z-score norm obtains the best performance in terms of image quality, it falls very short in terms of the expected estimation of the CA pattern, achieving suboptimal performance. Similar behavior is also obtained using the Min-Max norm, where pixel error is lower but diagnostic quality remains suboptimal. In both cases, the TI-PAN model shows a decrease in performance and exhibits a critical limitation in the clinical scenario, despite using TI-Loss. In the case of TI-norm, it guarantees the prevalence of the TI curve to be leveraged by the TI-Loss during training. These features allow obtaining the best scores in the estimation of the CA pattern while achieving similar image-based quality (25.4 in PSNR). This makes TI-norm an important component in generating high-quality diagnostic information in an image synthesis setting, which has not been reported before in the medical imaging synthesis field.
@@ 3.5 Quantitative and Qualitative evaluation of synthesized Images
An evaluation of the visual quality of all the generated images at the spatial feature level was performed. Figure 4 shows some generated sample results for images in the test set using all of the evaluated models together. Additionally, as images in the test set had at least one annotated ROI, they were also compared visually. Finally, the difference heatmaps provide compelling visual evidence of the pipeline’s efficacy. For 1.5T images, the generated samples exhibit enough realism in the visual domain for nearly all models, with only minor deviations observed in specific tissue textures and particular details. This is coherent with the previous results where a narrow margin of difference was observed in the visual domain. However, in case of the 3T images, a smaller difference between the real and generated images is evidenced when using the TILoss, especially in the annotated region where the difference in the brighter zone is lower. Besides, table 1 also reports the MAE, SSIM and PSNR for all models at full image and ROI leves, where no single model obtains the best scores consistently. In contrast, no large margin is evidenced among models, and slight differences are found. This result is in line with other cutting-edge studies that have shown good results in creating useful images from post-contrast sequences  [18, 11].